OpenStudy (anonymous):

Given the quadratic function f(x) = 25x2 – 100x + 100, find a value of x such that f(x) = 25.

OpenStudy (saifoo.khan):

25 = 25x2 – 100x + 100 Solve.

OpenStudy (anonymous):

25-100 =25x2-100x 25x2-100x+75 =0 divide with 25 x2-4x+3=0 (x-1)(x-3)=0 x=1, x=3

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

nice to do

OpenStudy (anonymous):

=)

OpenStudy (anonymous):

agree with saif